MASARYK UNIVERSITY

Petr Mejzlík, Ph.D.

Department of Information Technology

Faculty of Informatics Masaryk University Burešova 20 602 00 Brno Czech Republic Office: +42-5-4132-1237 ext. 382 Home: +42-5-4621-1890 Fax: +42-5-4121-2568 Internet: mejzlik@muni.cz

MASARYK UNIVERSITY

Petr Mejzlík, Ph.D.

Department of Information Technology

Faculty of Informatics Masaryk University Burešova 20 602 00 Brno Czech Republic Office: +42-5-4132-1237 ext. 382 Home: +42-5-4621-1890 Fax: +42-5-4121-2568 Internet: mejzlik@muni.cz

MASARYK UNIVERSITY

Petr Mejzlík, Ph.D.

Department of Information Technology

Faculty of Informatics Masaryk University Burešova 20 602 00 Brno Czech Republic Office: +42-5-4132-1237 ext. 382 Home: +42-5-4621-1890 Fax: +42-5-4121-2568 Internet: mejzlik@muni.cz

MASARYK UNIVERSITY

Petr Mejzlík, Ph.D.

Department of Information Technology

Faculty of Informatics Masaryk University Burešova 20 602 00 Brno Czech Republic Office: +42-5-4132-1237 ext. 382 Home: +42-5-4621-1890 Fax: +42-5-4121-2568 Internet: mejzlik@muni.cz

MASARYK UNIVERSITY

Petr Mejzlík, Ph.D.

Department of Information Technology

Faculty of Informatics Masaryk University Burešova 20 602 00 Brno Czech Republic Office: +42-5-4132-1237 ext. 382 Home: +42-5-4621-1890 Fax: +42-5-4121-2568 Internet: mejzlik@muni.cz

MASARYK UNIVERSITY

Petr Mejzlík, Ph.D.

Department of Information Technology

Faculty of Informatics Masaryk University Burešova 20 602 00 Brno Czech Republic Office: +42-5-4132-1237 ext. 382 Home: +42-5-4621-1890 Fax: +42-5-4121-2568 Internet: mejzlik@muni.cz

MASARYK UNIVERSITY

Petr Mejzlík, Ph.D.

Department of Information Technology

Faculty of Informatics Masaryk University Burešova 20 602 00 Brno Czech Republic Office: +42-5-4132-1237 ext. 382 Home: +42-5-4621-1890 Fax: +42-5-4121-2568 Internet: mejzlik@muni.cz

MASARYK UNIVERSITY

Petr Mejzlík, Ph.D.

Department of Information Technology

Faculty of Informatics Masaryk University Burešova 20 602 00 Brno Czech Republic Office: +42-5-4132-1237 ext. 382 Home: +42-5-4621-1890 Fax: +42-5-4121-2568 Internet: mejzlik@muni.cz

MASARYK UNIVERSITY

Petr Mejzlík, Ph.D.

Department of Information Technology

Faculty of Informatics Masaryk University Burešova 20 602 00 Brno Czech Republic Office: +42-5-4132-1237 ext. 382 Home: +42-5-4621-1890 Fax: +42-5-4121-2568 Internet: mejzlik@muni.cz

MASARYK UNIVERSITY

Petr Mejzlík, Ph.D.

Department of Information Technology

Faculty of Informatics Masaryk University Burešova 20 602 00 Brno Czech Republic Office: +42-5-4132-1237 ext. 382 Home: +42-5-4621-1890 Fax: +42-5-4121-2568 Internet: mejzlik@muni.cz

MASARYK UNIVERSITY

Petr Mejzlík, Ph.D.

Department of Information Technology

Faculty of Informatics Masaryk University Burešova 20 602 00 Brno Czech Republic Office: +42-5-4132-1237 ext. 382 Home: +42-5-4621-1890 Fax: +42-5-4121-2568 Internet: mejzlik@muni.cz

MASARYK UNIVERSITY

Petr Mejzlík, Ph.D.

Department of Information Technology

Faculty of Informatics Masaryk University Burešova 20 602 00 Brno Czech Republic Office: +42-5-4132-1237 ext. 382 Home: +42-5-4621-1890 Fax: +42-5-4121-2568 Internet: mejzlik@muni.cz